OpenStudy (anonymous):

fravorite type of movies action 180 drama 200 mystery 80 comedy 360 what is the ratio of the number of people who prefer mystery movies to drama

OpenStudy (susanna):

Ration --> mystery/drama

OpenStudy (anonymous):

9 to 10

OpenStudy (anonymous):

Divide 180 and 200 by 20. then divide your answer by 10. that is what i did.

OpenStudy (susanna):

80/200 8/20 2/5

OpenStudy (anonymous):

my mom is making a movie so i can help susann as right. i think

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

@ susanna I don't get the same answer as you when I do that all I get is 0.4

OpenStudy (susanna):

It's because 2/5=0.4

OpenStudy (susanna):

It's the answer, not the ratio.

OpenStudy (anonymous):

2/5 is the right answer I just don't know how to get there

OpenStudy (susanna):

You set up a ratio of mystery to drama. mystery/drama. Mystery --> 80 Drama --> 200 80/200 Then, simplify. 8/20 2/5

OpenStudy (mathstudent55):

A ratio is a comparison of two numbers by dividing one by the other. One way of showing a ratio is a s a fraction since a fraction means division. The ratio of A to B is the division of A by B. Write it as a fraction and reduce it.

OpenStudy (mathstudent55):

The ratio of people who prefer mystery movies to drama movies is found by looking up the two numbers in that order: number of people who prefer mystery movies: 80 number of people who prefer drama movies: 200 The ratio is: \(\dfrac{80}{200} \) Now just reduce the fraction: \(\dfrac{80}{200} = \dfrac{40 \times 2}{40 \times 5} = \dfrac{2}{5}\) BTW, this is just a more detailed explanation, but @Susanna did a good job above explaining it.